# Parsimonious Gaussian mixture models with piecewise-constant eigenvalue profiles

### v1

@@ 5.4 Clustering benchmark
[…]
Table 2: Comparison of the clustering ARIs (×$\times 100$) reached by the MPSA-H/R/U/D models at the end of the CPEM algorithm, the standard GMMs, and the HDDC and spectral clustering models, for various challenging data distributions. “--” indicates an average ARI of 00 or less, meaning that the clustering is similar or worse than random. The leading clustering ARIs are in bold.
@@ 6 Application to single-image denoising
Beyond traditional learning tasks like clustering, GMMs provide a flexible method for density modeling and therefore come with many other applications. In this section, we describe the application of MPSA models to single-image denoising.
@@ 6.1 Image denoising with GMMs
Many denoising methods rely on the self-similarity assumption [28, 9]. This regularity and periodicity prior implies that the patches extracted from images gather in some clusters [43]. Under this hypothesis, GMMs appear as natural models for the image patches. [70] run a GMM on a large database of natural image patches and use it for image denoising, deblurring and inpainting, outperforming all the traditional methods. However, this approach is not suitable for the task of single-image denoising, where we just have one image as an input, that is moreover noisy; hence the need for parsimonious generative models. Consequently, [36] propose to model the set of overlapping patches extracted from one image according to a low-dimensional MPPCA density with a shared mixture noise. The resulting model, HDMI (high-dimensional mixture models for unsupervised image denoising), automatically updates the intrinsic mixture dimensions $q_{c}$ during the EM based on a fixed hyperparameter $\sigma^{2}$ that models the variance of the image noise. More precisely, the intrinsic dimensions are chosen such that
@@
where the $\ell_{cj}$ are the eigenvalues of the local responsibility-weighted covariance matrices $S_{c}$. The noise level $\sigma^{2}$ can either be known (supervised denoising) or unknown (unsupervised denoising). In the latter case, it is estimated a posteriori as a hyperparameter, by minimizing the BIC of the learned mixture, similarly as in HDDC [11]. The next paragraph describes the extension of HDMI with MPSA models.
Consider an image $X\in\mathbb{R}^{s_{1}\times s_{2}}$ that has been corrupted by some isotropic Gaussian noise ${}{E\sim\mathcal{N}(0,\sigma^{2}_{\mathrm{true}}I_{s_{1}s_{2}})}$. Then the observed image can be written as $\tilde{X}=X+E$. We extract all the overlapping $s\times s$ patches from $\tilde{X}$. This yields a dataset $(\tilde{x}_{i})_{i=1}^{n}$ of noisy patches embedded in $\mathbb{R}^{p}$ with $p=s^{2}$. A nice illustration of such a dataset is provided in the Figure 1 of [36]. We assume that it follows an MPSA density (9), with the additional constraints $\lambda_{cd_{c}}=\sigma^{2}$, $\forall c$. We run Algorithm 2 to infer the mixture parameters $\hat{\theta}$. Proposition 5 then gives an explicit and efficient formula to denoise the patches.
@@
Let $\tilde{x}\in\mathbb{R}^{p}$ represent an $s\times s$ image patch following the noise model $\tilde{x}=x+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\sigma^{2}I_{p})$ is an isotropic Gaussian noise and $x\in\mathbb{R}^{p}$ is the original patch—sampled from an MPSA model of types $\hat{\gamma}$ and parameters $(\dots\hat{\pi}_{c},\hat{\mu}_{c},\hat{\Sigma}_{c}-\sigma^{2}I_{p}\dots)$. Then the denoised patch $\hat{x}\coloneqq\mathbb{E}(\,x\mid\tilde{x}\,)$ can be expressed as
@@
where $w_{c}=\mathbb{P}(\,y=c\mid\tilde{x}\,)$ (cf. Theorem 2).
Similarly as in Section 3.3, one can show that the use of the already-computed eigenvalues and eigenvectors and the substitution of the last eigenspace enables to switch from a computational complexity of $O(Cnp^{3})$ (with classical covariance inversion) to $O(Cnp^{2}q)$ (the most expensive step being the computation of ${}{\hat{\Pi}_{ck}=\hat{Q}_{ck}{\hat{Q}_{ck}}^{\top}}$). The denoised patches are finally aggregated by performing a uniform reprojection, yielding the denoised image $\hat{X}$.
@@ 6.2 Denoising experiments
We illustrate the previously presented approach on a $512\times 512$ image with (unknown) noise-deviation $\sigma_{\mathrm{true}}=30/255$ and patch size ${}{s=8}$. We fit the MPSA-H/R/U models with $C\in\{3,10,20\}$ and compare them to the full, spherical, and (supervised) HDMI mixture models. [^7] The HDMI models are learned in a supervised way (i.e. the true noise variance $\sigma^{2}_{\mathrm{true}}$ is given as an input), which is supposed to give a considerable advantage to HDMI over its competitors. We indeed found out that the results of HDMI were very sensitive with respect to the hyperparameter $\sigma^{2}$, and knowing that, we did not want to put arbitrary values in the grid search, which would have badly affected its performances. We evaluate the quality of the denoising using the partial signal-to-noise ratio evaluation metric ${}{\mathrm{PSNR}\coloneqq-10\log_{10}(\operatorname{MSE}(X,\hat{X}))}$, where MSE stands for mean squared error, similarly as in [36]. We repeat the experiment 5 times independently and report the results in Table 3.
@@
Finally, we propose in Figure 4 a heatmap-like visualization of the clusters directly on the image, for both (supervised) HDMI and MPSA-U models with $C=3$. More precisely, we map each patch to the cluster it most likely belongs to ($\hat{c}=\operatorname*{arg\,max}_{c\in[\![1,C]\!]}\mathbb{P}(\,y=c\mid\tilde{x}\,)$), and color this patch according to the number of parameters of the underlying mixture component (which is a PPCA model for HDMI and a PSA model for MPSA-U).
@@
This means that the uniform textures are attributed by HDMI an intrinsic dimension of respectively $9$ and $10$, while MPSA-U only requires respectively $5$ and $6$ dimensions. In contrast, the complex textures have an intrinsic dimension of $39$ with the HDMI model, while they have an intrinsic dimension of $20(=64-44)$ with the MPSA-U model.101010 A legitimate question might be: does an HDMI model of similar complexity as the MPSA-U model yield smaller PSNRs? The only way to reduce the number of parameters of the HDMI model being to increase the estimated noise (cf. Eq. (28)), we tried several values of $\sigma$. For $\sigma=40/255$, one gets the types $(1,63),(1,63),(1,63)$ and the PSNR $26.8$; for $\sigma=35/255$, one gets the types $(1,63),(1,63),(1^{2},62)$ and the PSNR $27.8$; for $\sigma=32/255$, one gets the types $(1,63),(1,63),(1^{7},57)$ and the PSNR $30.7$; for $\sigma=30/255(=\sigma_{\mathrm{true}})$, one gets the types $(1^{9},55),(1^{10},54),(1^{39},25)$ and the PSNR $31.1$. We notice an extreme sensitivity of the selected model complexity and PSNRs with respect to the choice of hyperparameter $\sigma$, but it seems that with similar numbers of parameters, HDMI and MPSA-U behave relatively similarly. Finally, we find that the heatmap obtained with MPSA-U is cleaner visually than the one of HDMI, probably due to the added parsimony.
@@ 7 Conclusion
We introduced a new family of parsimonious Gaussian mixture models with piecewise-constant covariance eigenvalue profiles. These were motivated by the important parsimony achieved by equalizing similar eigenvalues [59], going beyond classical low-rank Gaussian mixture models [61, 11]. While the unsupervised learning of the mixture parameters was easily enabled via an EM algorithm, the automatic inference of the groups of eigenvalues to equalize was more technical. It was finally achieved via a provably-monotonous componentwise penalized EM algorithm. The numerical experiments showed the superior likelihood–parsimony tradeoffs achieved by our models when compared to the two extreme cases: the full GMM (with distinct covariance eigenvalues) and the spherical GMM (with equal covariance eigenvalues), and the effective application of our models to clustering, density modeling and denoising tasks.
@@
Beyond our extensive experiments to analyze the learning algorithm, one could investigate more data regimes, variants and applications. We notably believe that our new CPEM algorithm has the potential to be applied to many classical low-rank mixture models, removing the need for post-training hyperparameter selection. Instead of an EM algorithm, one could also directly optimize the penalized log-likelihood with gradient descent techniques, drawing from the recent relaxation of parsimony of [58]. The MPSA models could also be naturally applied to the numerous generative modeling tasks for which classical GMMs are often used. Their expressivity could be improved by drawing from the recent advances in GMMs and machine learning [55, 45, 20].
@@
Other potential applications of our parsimonious models are manifold learning and surface reconstruction. Indeed, it is often assumed that the observed data has been sampled from a low-dimensional manifold (up to isotropic Gaussian noise). The density could be fit with our MPSA models, and the underlying shape could be recovered with classical surface reconstruction methods. The types of the mixture components could tell something about the local intrinsic dimensions of the underlying shape [1, 17].
@@
Finally, since the PSA models have been originally introduced to improve the interpretation of principal components with close eigenvalues, one should definitely investigate the interpretation of the principal subspaces found by the MPSA models. We conclude this paper with a simple example on image patches. As discussed and experimented in Section 6, image patches are classically represented with Gaussian mixtures, and we believe that they carry some invariances that could be well modeled with piecewise-constant eigenvalue profiles. Classical works indeed reveal some symmetries in the eigenmodes (cf. [52, FIG. 1], [46, Figure 5(a)] and [70, Figure 6]) that are reminiscent of the curse of isotropy of [59]. We illustrate this perspective by fitting an MPSA-U model with $C=3$ on patches extracted from the clean Simpson image. The output model types are quite interesting, e.g. $\gamma=(2,3,4,5,\dots)$ for the first cluster. We display in Figure 5 the first few principal components associated with the first cluster, gathered according to their eigenvalue multiplicities, and then sample some points from the associated principal subspaces. This essentially boils down to generating convex combinations of the principal components with similar eigenvalues.
@@
Figure 5: Illustration of the principal subspaces found by fitting an MPSA-U model to patches extracted from the clean Simpson image. The first cluster has the type $\gamma=(2,3,4,5,\dots)$, which leads to principal subspaces #$\#1,\#2,\#3,\#4$ of respective dimension $2$, $3$, $4$ and $5$. We display some samples from these principal subspaces (i.e. convex combinations of the associated principal components). These can be interpreted as feature subspaces of increasing frequency, with rotational invariance in the first subspace.
@@
We observe some interesting patterns within the principal subspaces, which can somewhat be interpreted as low-frequency feature subspaces with rotational invariance.

### v2

@@ 5.5 Clustering benchmark
We now perform some clustering experiments, where the learning procedure is similar to the previous experiments, but the reported metric is the adjusted rand index (ARI) [rand_objective_1971, hubert_comparing_1985], which is the classical substitute for the accuracy in clustering problems. This scores varies between (approximately) 0 for a random clustering and (exactly) $1$ for a perfect clustering. We add two competitors: HDDC [bouveyron_high-dimensional_2007] and spectral clustering [ng_spectral_2001]. The first three datasets are synthetic. They all have in common that $C=3$, $(\pi_{1},\pi_{2},\pi_{3})=(0.4,0.3,0.3)$, $(\lambda_{11},\lambda_{21},\lambda_{31})=(10,1,.1)$, $\lambda_{cp}/\lambda_{c1}=0.01\,\forall c$. Now, more specifically, the first dataset (MPSA $10$) has $n=200$, $p=10$, $\mu_{c}\sim\mathcal{U}\left(\left[-1,1\right]^{p}\right)$, and $\gamma_{1},\gamma_{2},\gamma_{3}=(1,p-1),(1,2,p-3),(1,2,4,p-7)$}. The second dataset (MPSA $50$) has the same characteristics as the previous one but with $\mu_{c}=0$ (mixed clusters, which is the case of the next ones), $p=50$ and $n=1000$. The third one (Full $10$) has $p=10$, $n=200$ and $(\gamma_{1},\gamma_{2},\gamma_{3})=((1^{p}),(1^{p}),(1^{p}))$}. [^7] Note that we choose particularly challenging settings (notably with close—or even exactly similar—centers) to show how our models can adapt the covariance structures, rather than the centers. We then have three real datasets (Ionosphere [sigillito_ionosphere_1989], Wine [aeberhard_wine_1991], Breast [wolberg_breast_1995]), respectively with $(n,p,C)=(351,34,2),(178,13,3),(569,30,2)$. The average clustering ARIs after 10 independent $n$-samples (or after a stratified $10$-fold cross-validation for the real datasets) are reported in Table 2.
Table 2: Comparison of the clustering ARIs ($\times 100$) reached by the MPSA-H/R/U/D models at the end of the CPEM algorithm, the standard GMMs, and the HDDC and spectral clustering models, for various challenging data distributions. “-” indicates an average ARI of 0 or less, meaning that the clustering is similar or worse than random. The leading clustering ARIs are in bold.
@@
5.6 Running times
As seen in the previous experiments, the four strategies are rather comparable in terms of likelihood–parsimony tradeoffs, density estimation and clustering, with the MPSA-R strategy being overall slightly worse than the other strategies. This subsection aims to compare the different strategies from the running time perspective.
We generate data from the MPSA generative model, with $n=1000$, $C=3$, $(\gamma_{1},\gamma_{2},\gamma_{3})=((1,p-1),(1,2,p-3),(1,2,4,p-7))$, $(\pi_{1},\pi_{2},\pi_{3})=(0.4,0.3,0.3)$, $\mu_{c}=5$, $(\lambda_{11},\lambda_{21},\lambda_{31})=(3,2,1)$, and $\lambda_{cp}/\lambda_{c1}=0.01\,\forall c$. The experiments are repeated independently 10 times. We display the running times of the different GMMs as a function of the dimension $p$ in Figure 5.
Figure 5: Running time benchmark for increasing dimension.
The spherical GMM is significantly faster than all the other models—which makes sense since the estimation of the isotropic covariance matrices does not involve any matrix inversion or eigenvalue decomposition. The full GMM is faster than the MPSA models, in average, but it scales like the MPSA-R and MPSA-U strategies. The MPSA-H and MPSA-D strategies, for their part, are much slower, making them less adapted in high dimensions ($p\gtrapprox 1000$ for MPSA-H and $p\gtrapprox 100$ for MPSA-D). Let us now compare the four MPSA strategies.
MPSA-R is significantly faster than the other strategies—at the cost of apparently lower modeling performances, as we could see in the previous experiments. MPSA-U is the second fastest strategy, scaling like MPSA-R, although possibly requiring more iterations than MPSA-H, as seen in Figure 2. MPSA-H is the third, as it involves a hierarchical clustering of the eigenvalues, which can become pretty costly in high dimensions (cf. Section 2). MPSA-D is the slowest—by far—as it starts from the full covariance model, and decreases the number of parameters gradually, which can take a lot of iterations to converge if the true model is parsimonious (cf. Figure 2). Hence, the empirical running times rather match the previously built intuitions.
The comparisons between the MPSA and GMMs are to nuance, as our code is certainly not as much optimized as the one of scikit-learn [pedregosa_scikit-learn_2011], from which the classical GMMs are drawn. Moreover, the current version of the code does not include certain numerical tricks such as the previously evoked truncated SVD, which is investigated in Section B.4. We could expect well-engineered MPSA models to be significantly faster than full GMMs, as truncated SVDs are significantly faster than covariance inversion.
As a conclusion to the series of experiments performed in this section, we would probably recommend to adopt the MPSA-U strategy by default in general settings. Indeed, it is the second fastest strategy (scaling as the fastest one and being much faster than the third and fourth), and it has a good control on parsimony (as it is initialized with isotropic covariances and it increases gradually the number of parameters throughout the CPEM iterations, by finding the most informative gaps in the eigenvalue profiles). In contrast, the other strategies can possibly select overparameterized models in the first iterations, which can negatively impact learning in high dimensions.
@@ 6 Application to single-image denoising
Beyond traditional learning tasks like clustering, GMMs provide a flexible method for density modeling and therefore come with many other applications. In this section, we briefly describe the application of MPSA models to single-image denoising. More details about the denoising framework can be found in houdard_high-dimensional_2018.
@@ 6.1 Image denoising with GMMs
Many denoising methods rely on the self-similarity assumption: any small patch extracted from an image has “similar” patches in the same image [efros_texture_1999, buades_non-local_2005]. This regularity and periodicity prior implies that patches extracted from natural images form well-defined clusters (see [lebrun_secrets_2012] for more details). Such a redundancy enables to reduce the noise variance by averaging the similar patches. Under this self-similarity hypothesis, GMMs appear as natural models for the set of patches extracted from one image.
The celebrated work of [zoran_learning_2011] runs a GMM on a large database of natural image patches and use it for image denoising, deblurring and inpainting, outperforming all the traditional methods. However, this approach is not suitable for the task of single-image denoising, where we just have one image as an input, that is moreover noisy; hence the need for parsimonious generative models. Consequently, [houdard_high-dimensional_2018] propose to model the set of overlapping patches extracted from one image according to a low-dimensional MPPCA density with a shared mixture noise. The resulting model, HDMI (high-dimensional mixture models for unsupervised image denoising), automatically updates the intrinsic mixture dimensions $q_{c}$ during the EM based on a fixed hyperparameter $\sigma^{2}$ that models the variance of the image noise. More precisely, the intrinsic dimensions are chosen such that
@@
where the $\ell_{cj}$ are the eigenvalues of the local responsibility-weighted covariance matrices $S_{c}$. The noise level $\sigma^{2}$ can either be known (supervised denoising) or unknown (unsupervised denoising). In the latter case, it is estimated a posteriori as a hyperparameter, by minimizing the BIC of the learned mixture, similarly to HDDC [bouveyron_high-dimensional_2007]. The next paragraph describes the extension of HDMI with MPSA models.
Consider an image $X\in\mathbb{R}^{s_{1}\times s_{2}}$ that has been corrupted by some isotropic Gaussian noise $E\sim\mathcal{N}(0,\sigma^{2}_{\mathrm{true}}I_{s_{1}s_{2}})$}. Then the observed image can be written as $\tilde{X}=X+E$. We extract all the overlapping $s\times s$ patches from $\tilde{X}$, from left to right, top to bottom. This yields a dataset $(\tilde{x}_{i})_{i=1}^{n}$ of $n=(s_{1}-s+1)(s_{2}-s+1)$ noisy patches embedded in $\mathbb{R}^{p}$ with $p=s^{2}$. A nice illustration of such a dataset is provided in the Figure 1 of [houdard_high-dimensional_2018]. We assume that it follows an MPSA density (9), with the additional constraints $\lambda_{cd_{c}}=\sigma^{2}$, $\forall c$. We run Algorithm 2 to infer the mixture parameters $\hat{\theta}$. Proposition 5 then gives an explicit and efficient formula to denoise the patches.
@@
[…]
@@
[…]
Similarly to Section 3.3, one can show that the use of the already-computed eigenvalues and eigenvectors and the substitution of the last eigenspace enables to switch from a computational complexity of $O(Cnp^{3})$ (with classical covariance inversion) to $O(Cnp^{2}q)$ (the most expensive step being the computation of $\hat{\Pi}_{ck}=\hat{Q}_{ck}{\hat{Q}_{ck}}^{\top}$}). The denoised patches are finally aggregated by performing a uniform reprojection—i.e., positioning back the denoised patches in the image from left to right, top to bottom and averaging them in the overlapping regions—yielding the denoised image $\hat{X}$.
@@ 6.2 Denoising experiments
We illustrate the previously presented approach on a $512\times 512$ image with (unknown) noise-deviation $\sigma_{\mathrm{true}}=30/255$ and patch size s=8~{s=8}. We fit the MPSA-H/R/U models with $C\in\{3,10,20\}$ and compare them to the full, spherical, and (supervised) HDMI mixture models. [^8] The HDMI models are learned in a supervised way (i.e. the true noise variance $\sigma^{2}_{\mathrm{true}}$ is given as an input), which is supposed to give a considerable advantage to HDMI over its competitors. We indeed found out that the results of HDMI were very sensitive with respect to the hyperparameter $\sigma^{2}$, and knowing that, we did not want to put arbitrary values in the grid search, which would have badly affected its performances. We evaluate the quality of the denoising using the partial signal-to-noise ratio evaluation metric ${\mathrm{PSNR}\coloneqq-10\log_{10}(\operatorname{MSE}(X,\hat{X}))}$, where MSE stands for mean squared error, similarly to [houdard_high-dimensional_2018]. We repeat the experiment 5 times independently and report the results in Table 3.
@@
[…]
@@
This means that the uniform textures are attributed by HDMI an intrinsic dimension of respectively $9$ and $10$, while MPSA-U only requires respectively $5$ and $6$ dimensions. In contrast, the complex textures have an intrinsic dimension of $39$ with the HDMI model, while they have an intrinsic dimension of 20(=6444)20(=64-44) with the MPSA-U model. [^11] A legitimate question might be: does an HDMI model of similar complexity as the MPSA-U model yield smaller PSNRs? The only way to reduce the number of parameters of the HDMI model being to increase the estimated noise (cf. Eq. (28)), we tried several values of $\sigma$. For $\sigma=40/255$, one gets the types $(1,63),(1,63),(1,63)$ and the PSNR $26.8$; for $\sigma=35/255$, one gets the types $(1,63),(1,63),(1^{2},62)$ and the PSNR $27.8$; for $\sigma=32/255$, one gets the types $(1,63),(1,63),(1^{7},57)$ and the PSNR $30.7$; for $\sigma=30/255(=\sigma_{\mathrm{true}})$, one gets the types $(1^{9},55),(1^{10},54),(1^{39},25)$ and the PSNR $31.1$. We notice an extreme sensitivity of the selected model complexity and PSNRs with respect to the choice of hyperparameter $\sigma$, but it seems that with similar numbers of parameters, HDMI and MPSA-U behave relatively similarly. Finally, we find that the heatmap obtained with MPSA-U is cleaner visually than the one of HDMI, probably due to the added parsimony.
Remark 11 (Influence of $s$ and $C$ on denoising).
In all the experiments of this section, the patch size is set to $s=8$ and the number of components $C$ varies heuristically according to the denoising task. We discuss further in depth the influence of $s$ and $C$ on the denoising performances and running times in Section B.3. Similar experiments can be found in [houdard_high-dimensional_2018].
@@ 7 Conclusion
We introduced a new family of parsimonious Gaussian mixture models with piecewise-constant covariance eigenvalue profiles. These were motivated by the important parsimony achieved by equalizing similar eigenvalues [szwagier_curse_2025], going beyond classical low-rank Gaussian mixture models [tipping_mixtures_1999, bouveyron_high-dimensional_2007]. While the unsupervised learning of the mixture parameters was easily enabled via an EM algorithm, the automatic inference of the groups of eigenvalues to equalize was more technical. It was finally achieved via a provably-monotonous componentwise penalized EM algorithm. The numerical experiments showed the superior likelihood–parsimony tradeoffs achieved by our models when compared to the two extreme cases: the full GMM (with distinct covariance eigenvalues) and the spherical GMM (with equal covariance eigenvalues), and the effective application of our models to clustering, density modeling and denoising tasks.
@@
Beyond our extensive experiments to analyze the learning algorithm, one could investigate more data regimes, variants and applications. We notably believe that our new CPEM algorithm has the potential to be applied to many classical low-rank mixture models, removing the need for post-training hyperparameter selection. Instead of an EM algorithm, one could also directly optimize the penalized log-likelihood with gradient descent techniques [hosseini_alternative_2020], drawing from the recent relaxation of parsimony of [szwagier_eigengap_2025]. The MPSA models could also be naturally applied to the numerous generative modeling tasks for which classical GMMs are often used. Their expressivity could be improved by drawing from the recent advances in GMMs and machine learning [richardson_gans_2018, loconte_subtractive_2023, byrski_gaussian_2024]. Robustness could be achieved by integrating our principles into the recent mixtures of high-dimensional elliptical distributions of oudoumanessah_scalable_2024.
@@
Other potential applications of our parsimonious models are manifold learning and surface reconstruction. Indeed, it is often assumed that the observed data has been sampled from a low-dimensional manifold (up to isotropic Gaussian noise). The density could be fit with our MPSA models, and the underlying shape could be recovered with classical surface reconstruction methods. The types of the mixture components could tell something about the local intrinsic dimensions of the underlying shape [akhoj_principal_2023, buet_flagfolds_2023].
@@
Finally, since the PSA models have been originally introduced to improve the interpretation of principal components with close eigenvalues, one should definitely investigate the interpretation of the principal subspaces found by the MPSA models. We conclude this paper with a simple example on image patches. As discussed and experimented in Section 6, image patches are classically represented with Gaussian mixtures, and we believe that they carry some invariances that could be well modeled with piecewise-constant eigenvalue profiles. Indeed, many classical works have noticed that the principal components in natural image patch datasets have repeated eigenvalues, accounting for shared frequencies (see, for instance, olshausen_emergence_1996, mairal_online_2010 and zoran_learning_2011). Therefore, the MPSA models could reveal these hidden symmetries by detecting the multiple covariance eigenvalues in the mixture components, similarly to [szwagier_curse_2025]. We illustrate this perspective by fitting an MPSA-U model with $C=3$ on patches extracted from the clean Simpson image. The output model types are quite interesting, e.g. $\gamma=(2,3,4,5,\dots)$ for the first cluster. We display in Figure 8 the first few principal components associated with the first cluster, gathered according to their eigenvalue multiplicities, and then sample some points from the associated principal subspaces. This essentially boils down to generating convex combinations of the principal components with similar eigenvalues.
@@
Figure 8: Illustration of the principal subspaces found by fitting an MPSA-U model to patches extracted from the clean Simpson image. The inferred eigenvalue multiplicities for the first component are $\gamma_{1}=(2,3,4,5,\dots)$, which means that the principal subspaces #1,#2,#3,#4\#1,\#2,\#3,\#4 have respective dimensions 22, 33, 44 and $5$. We display some samples from these principal subspaces (more precisely, convex combinations of the associated principal components). These can be interpreted as feature subspaces with (limited) frequency invariance, which are reminiscent of the findings of [szwagier_curse_2025] about the curse of isotropy.
@@
We observe some interesting patterns within the principal subspaces, which can somewhat be interpreted as feature subspaces with (limited) frequency invariance.